OpenStudy (anonymous):

pls what is the limit as x turns to 0 of tan3x/sin8x

OpenStudy (ddcamp):

Use L'hospitals rule.

OpenStudy (anonymous):

is it 3/64

OpenStudy (ddcamp):

I don't know. Using L'hospitals rule, we have: \[\lim_{x \rightarrow 0}\frac{ \tan(3x) }{ \sin(8x) } = \lim_{x \rightarrow 0}\frac{ 3\sec^2(3x) }{ 8\cos(8x) }\] Can you find a value for the second expression?

OpenStudy (anonymous):

its true the answer is 3/8 by leting x=0